# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v1

@@ B.1 Datasets
We cover our generated synthetic datasets in Section 4, vision datasets in Section 5, text datasets in Section 6, and the datasets for robustness in Section 7.
@@ B.1.1 Synthetic Datasets
We generate two synthetic datasets to validate the correctness of our IP method in measuring the convex and non-convex models’ utility, fairness, and robustness. Specifically, we generate a linear dataset using the scikit-learn library’s make_blobs function, [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_blobs.html which consists of 150 training samples and 100 test samples. The second dataset is the non-linear half moons dataset so that we can train a Multi-Layer Perception network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. [^3] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html Here we manually flip the labels of 20 samples (10 from each class) to add noise to the data.
@@ B.1.2 Vision Datasets
Both the CIFAR-10N and CIFAR-100N datasets [59] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [48], respectively. Each input is a 32x32 RGB image with a dimension of (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50000 image-label pairs, and the test set contains 10000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst, which is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate, which is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random which has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. For CIFAR-100N there is only a single noisy setting (Noisy100) due to the large number of labeling classes, and the overall noise rate is 40.20%.
@@ B.1.3 NLP Datasets
For the three GLUE datasets—RTE, CoLA, and QNLI[51], we perturbed the validation set as follows. The Recognizing Textual Entailment (RTE) dataset consists of sentence pairs labeled as entailment or not entailment. It is derived from a series of annual textual entailment challenges. The dataset includes 2,490 training examples and 277 validation examples. Since the test set does not have labels, we split the validation set into two parts, with one half used as the validation set for computing influence, and the other half used as the test set. Subsequently, we generated perturbed datasets for both the validation set and the test set using a seq2seq model [53]. The Corpus of Linguistic Acceptability (CoLA) dataset consists of sentences labeled as grammatically acceptable or unacceptable. This dataset is derived from publications in linguistics and includes 8,551 training examples and 1,043 validation examples. As with RTE, the test set for CoLA does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were generated using a seq2seq model [53] for both the validation set and the test set. The Question Natural Language Inference (QNLI) dataset is a large-scale corpus for question answering, consisting of question-sentence pairs from the Stanford Question Answering Dataset (SQuAD). The task is to determine whether the context sentence contains the answer to the question. It includes 104,743 training examples and 5,463 validation examples. Similar to the other datasets, the test set does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were also generated using a seq2seq model [53] for both the validation set and the test set. These datasets represent diverse natural language understanding tasks, and using them for fine-tuning helps evaluate the fairness and utility of the fine-tuned RoBERTa model.
@@ B.1.4 Datasets for Defencing Attacks
We utilize three datasets: Bank, CelebA, and JigsawToxicity, to evaluate the defense against adaptive evasion adversaries using a Logistic Regression model. The Bank dataset [56] consists of features extracted from direct marketing campaigns of a Portuguese banking institution. The goal is to predict whether a client will subscribe to a term deposit. The dataset includes 18,292 training examples, 6,098 validation examples, and 6,098 test examples. The feature dimension of the dataset is 51. Both the validation and test sets have labels. The CelebA dataset [57] is a large-scale face attributes dataset with more than 200,000 celebrity images, each annotated with 40 binary attributes. For this experiment, we focus on a subset of these images. The dataset is split into 62,497 training examples, 20,833 validation examples, and 20,833 test examples. The feature dimension of the dataset is 39. Both the validation and test sets have labels. The JigsawToxicity dataset [58] contains a large number of comments from Wikipedia labeled for toxicity. The goal is to predict the toxicity level of a given comment. The dataset includes 18,000 training examples, 6,000 validation examples, and 6,000 test examples. The feature dimension of the dataset is 385. These datasets cover a range of tasks from marketing prediction and face attribute recognition to toxicity detection, providing a robust evaluation of the model’s defense mechanisms against adaptive adversaries.
@@ B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [50] architecture used as the base model for the noisy vision datasets, then the RoBERTa [52] NLP transformer model. We also describe implementation details and parameter values for the label correction baselines used in Sections 7 the influence-based baselines used throughout the paper.
@@ B.2.1 MLP
In our experiments (as described in Section 4), we utilized a three-layer Multi-Layer Perceptron (MLP) classifier. This MLP is structured with three fully connected layers, each followed by a ReLU activation function. Specifically, the input layer maps the input features to 32 neurons, the hidden layer maintains this dimensionality, and the output layer is a single neuron producing the final prediction. The final layer’s output is passed through a sigmoid activation function to yield a probability score.
@@ B.2.3 RoBERTa
We use the Roberta-base model from Huggingface [^4] https://huggingface.co/docs/transformers/model_doc/roberta in Section 6. A learning rate of 0.00001 and the batch sizes of RTE, CoLA, and QNLI are 64, 16, and 32. The model is fine-tuned over 10 epochs. The loss function used is a negative log-likelihood as the datasets are all for binary classification.
@@ B.2.4 Logistic Regression
The Logistic Regression model in Section 7 is implemented using the sklearn library. The model uses L2 regularization with and a maximum iteration limit of 2048 [2].
@@ B.2.5 Influence-Based Baselines
In our experiments, we utilize the following influence function-based methods as baselines: TracIn [16] replaces the Hessian matrix with the identity matrix; LiSSA [10] and EKFAC [13] use implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [3] swaps the order of matrix multiplication to obtain a closed-form expression; Self-TracIn [18] and Self-LiSSA [17] are self-expression versions of TracIn and LiSSA, where $\nabla v^{\textrm{util}}$ is replaced with $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$. Our IP closely resembles TracIn, with the added application of dropout with $\mathcal{U}(0,0.01)$ on the model parameters. Additionally, our IP Ensemble constitutes an ensemble version of IP with an ensemble size of 5.
@@ C.1 IP for measuring fairness and robustness
Similar to Figure 2, we conduct the experiments based on Eqs. (3) and (4) to demonstrate the effectiveness of our IP in measuring the sample influence on fairness and robustness. Figure 8 shows the relationship between IP and the traditional influence function, indicating IP is a good surrogate of the traditional influence function in measuring fairness and robustness as well.
Figure 8: Illustrating our inner product approach on measuring fairness and robustness. A and D illustrate the same 2D linearly separable synthetic dataset in Figure 2 trained using a Logistic Regression model for binary classification, where the solid and dashed point boundaries denote the majority and minority subgroups. B and E represent the validation set. C and F show the estimated influence on fairness and robustness by the traditional influence function and our IP method.
@@ C.2 Performance of IP Ensemble with different rates of removed samples
In our paper, we explore different rates of removed samples on CIFAR-10N-a. Table 4 shows the performance of IP Ensemble with different rates of removed samples. Except for not removing, there is no significant difference in the remaining results.
@@ C.3 Parameter Analysis on Dropout Rate and Ensemble Size
C.4 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Table 6 shows the time complexity of various influence function-based methods. Except for the vanilla calculation, all other methods have linear time complexity in terms of the sample size. However, they have large divergence in real execution time.
Table 6: Computational complexity of various influence-function-based methods ($n$ is #training samples and $p$ is #model parameters, $k$ is #ensemble size. "-" denotes no runs.)
@@ C.4 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Self-TracIn [18] Self-influence $\mathcal{O}(np)$ 0.15 1.41 0.15
@@ Appendix D Broader Impact and Limitations
Our work aims to address issues that currently hinder the applicability of influence estimation in deep learning models. By enabling influence estimation for deep models, practitioners can assess whether training samples are beneficial or detrimental to performance, making models more interpretable and performant. As we demonstrate through extensive experiments across multiple problem settings, our proposed outlier gradient analysis approach outperforms existing baselines and can augment model performance by trimming detrimental samples. As a result, our work paves the way for significant positive societal impact, especially with the increased adoption of larger and deeper neural networks such as LLMs. However, despite our research demonstrating superior performance and potential benefits, there are several limitations to consider. First, the simplicity of our Hessian-free approximation method, while effective in many scenarios, may not capture the full complexity of more nuanced model behaviors and interactions within the data. Additionally, our method assumes a level of model and data homogeneity that might not hold in highly heterogeneous datasets or more complex model architectures. Furthermore, while our IP Ensemble strategy enhances generalization by simulating diverse models, it introduces additional computational overhead and complexity that may not be feasible for all applications or environments, particularly those with limited computational resources. There is also the consideration of robustness and fairness across different domains and types of data; our current evaluations, while extensive, may not cover all possible scenarios or edge cases, potentially limiting the generalizability of our findings. Finally, ethical considerations around data trimming and sample selection must be carefully managed to avoid unintended biases or negative impacts on model performance for underrepresented groups. As influence estimation becomes more integrated into model training and evaluation pipelines, ongoing research and monitoring will be essential to ensure that these techniques are applied responsibly and equitably.

### v2

@@ B.1 Datasets
We cover our generated synthetic datasets in Section 4, vision datasets with noisy labels in Section 5, text datasets in Section 6, and datasets for robustness in Section 7.
@@ B.1.1 Synthetic Datasets
We generate two synthetic datasets to validate the correctness of our IP method in measuring the convex and non-convex models’ utility, fairness, and robustness. Specifically, we generate a linear dataset using the scikit-learn [45] library’s make_blobs function, which consists of 150 training samples and 100 test samples. The second dataset is the non-linear half moons dataset so that we can train a Multi-Layer Perception network with two hidden layers with ReLU activations. The training set has 250 samples and the test set has 100 samples, and the dataset is generated using the scikit-learn library’s make_moons function. Here we manually flip the labels of 10 samples from each class to add noise to the data.
@@ B.1.2 Vision Datasets
Both the CIFAR-10N and CIFAR-100N datasets [43] consist of the same input images that make up the CIFAR-10 (10 classes) and CIFAR-100 (100 classes) datasets [29], respectively. Each input is a 32x32 RGB image with a dimension of (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels are noisy, as they contain real-world human annotation errors collected using 3 annotators on Amazon Mechanical Turk. As these datasets are based on human-annotated noise, they model noisy real-world datasets more realistically, compared to synthetic data alternatives. The training set for both datasets contains 50,000 image-label pairs, and the test set contains 10,000 image-label pairs that are free from noise. For CIFAR-10N we utilize three noise settings for experiments in the paper– (1) Worst is the dataset version with the highest noise rate (40.21%) as the worst possible annotation label for the image is chosen, (2) Aggregate is the least noisy dataset (9.03%) as labels are chosen via majority voting amongst the annotations, and (3) Random has intermediate noise (17.23%) and consists of picking one of the annotators’ labels. We use the first annotator for the random labels. Due to the substantial number of labeling classes in CIFAR-100N, only a single noisy setting is available, with an overall noise rate of 40.20%.
@@ B.1.3 NLP Datasets
For the three GLUE datasets—RTE, CoLA, and QNLI[56], we perturbe the validation set as follows. The Recognizing Textual Entailment (RTE) dataset consists of sentence pairs labeled as entailment or not entailment. It is derived from a series of annual textual entailment challenges. The dataset includes 2,490 training examples and 277 validation examples. Since the test set does not have labels, we split the validation set into two parts, with one half used as the validation set for computing influence, and the other half used as the test set. Subsequently, we generated perturbed datasets for both the validation set and the test set using a seq2seq model [47]. The Corpus of Linguistic Acceptability (CoLA) dataset consists of sentences labeled as grammatically acceptable or unacceptable. This dataset is derived from publications in linguistics and includes 8,551 training examples and 1,043 validation examples. As with RTE, the test set for CoLA does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were generated using a seq2seq model [47] for both the validation set and the test set. The Question Natural Language Inference (QNLI) dataset is a large-scale corpus for question answering, consisting of question-sentence pairs from the Stanford Question Answering dataset. The task is to determine whether the context sentence contains the answer to the question. It includes 104,743 training examples and 5,463 validation examples. Similar to the other datasets, the test set does not have labels, so we split the validation set into two parts, using one half as the validation set for computing influence, and the other half as the test set. Perturbed datasets were also generated using a seq2seq model [47] for both the validation set and the test set. These datasets represent diverse natural language understanding tasks, and using them helps evaluate the fairness and utility of the fine-tuned RoBERTa model.
@@ B.1.4 Datasets for Defencing Attacks
We utilize three datasets: Bank, CelebA, and JigsawToxicity, to evaluate the defense against adaptive evasion adversaries. The Bank dataset [40] consists of features extracted from direct marketing campaigns of a Portuguese banking institution. The goal is to predict whether a client will subscribe to a term deposit. The dataset includes 18,292 training examples, 6,098 validation examples, and 6,098 test examples. The feature dimension of the dataset is 51. Both the validation and test sets have labels. The CelebA dataset [37] is a large-scale face attributes dataset with more than 200,000 celebrity images, each annotated with 40 binary attributes. For this experiment, we focus on a subset of these images. The dataset is split into 62,497 training examples, 20,833 validation examples, and 20,833 test examples. The feature dimension of the dataset is 39. Both the validation and test sets have labels. The JigsawToxicity dataset [41] contains a large number of comments from Wikipedia labeled for toxicity. The goal is to predict the toxicity level of a given comment. The dataset includes 18,000 training examples, 6,000 validation examples, and 6,000 test examples. The feature dimension of the dataset is 385. These datasets cover a range of tasks from marketing prediction and face attribute recognition to toxicity detection, providing a robust evaluation of the model’s defense mechanisms against adaptive adversaries.
@@ B.2 Models and Methods
We now describe the models and the methods used in our experiments throughout the main paper. First, we describe the ResNet-34 [22] architecture used as the base model for the noisy vision datasets, then the RoBERTa [35] NLP transformer model. We also describe implementation details and parameter values for the label correction baselines used in Sections 7 and the influence-based baselines used throughout the paper.
@@ B.2.1 MLP
In our experiments as described in Section 4, we utilized a three-layer Multi-Layer Perceptron (MLP) classifier. This MLP is structured with three fully connected layers, each followed by a ReLU activation function. Specifically, the input layer maps the input features to 32 neurons, the hidden layer maintains this dimensionality, and the output layer is a single neuron producing the final prediction. The final layer’s output is passed through a sigmoid activation function, generating a probability score for prediction.
@@ B.2.3 RoBERTa
We use the Roberta-base model from Huggingface [^2] https://huggingface.co/docs/transformers/model_doc/roberta in Section 6. The learning rate is 0.00001 and the batch sizes of RTE, CoLA, and QNLI are 64, 16, and 32. The model is fine-tuned over 10 epochs. The loss function used is a negative log-likelihood as the datasets are all for binary classification.
@@ B.2.4 Logistic Regression
The Logistic Regression model in Section 7 is implemented using the sklearn library. The model uses L2 regularization with and a maximum iteration limit of 2,048 [11].
@@ B.2.5 Influence-Based Baselines
In our experiments, we utilize the following influence function-based methods as baselines: TracIn [46] replaces the Hessian matrix with the identity matrix; LiSSA [26] and EKFAC [19] use implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [31] swaps the order of matrix multiplication to obtain a closed-form expression; Self-TracIn [52] and Self-LiSSA [5] are self-expression versions of TracIn and LiSSA, where $\nabla v^{\textrm{util}}$ is replaced with $\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta})$. GEX [25] utilizes ensemble methods based on checkpoints from extra stochastic gradient descent on the converged model. TDA [3] focuses checkpoints during the training process, ensembles the influence via EKFAC [19]. Our IP closely resembles TracIn and GEX, with the added application of dropout with $\mathcal{U}(0,0.01)$ on the model parameters. TracIn, GEX, IP, and IP Ensemble are all Hessian-free and they are ensemble versions, except for IP. Here, we present their calculations to highlight the differences between them as follows:
$\begin{split}\mathcal{I}^{\textrm{util}}_{\textup{IP}}(-z_{i},\hat{\theta})&=\sum\nolimits_{z_{j}\in V}\nabla_{\hat{\theta}}\ell(z_{j};\hat{\theta})^{\top}\nabla_{\hat{\theta}}\ell(z_{i};\hat{\theta}),\\ \mathcal{I}_{\textrm{TracIn}}^{\textrm{util}}(-z_{i},\Theta_{\textrm{TracIn}})&=\frac{1}{T}\sum\nolimits_{\hat{\theta}_{t}\in\Theta_{\textrm{TracIn}}}\mathcal{I}^{\textrm{util}}_{\textup{IP}}(-z_{i},\hat{\theta}_{t}),\\ \mathcal{I}_{\textrm{GEX}}^{\textrm{util}}(-z_{i},\Theta_{\textrm{GEX}})&=\frac{1}{T}\sum\nolimits_{\hat{\theta}_{t}\in\Theta_{\textrm{GEX}}}\mathcal{I}^{\textrm{util}}_{\textup{IP}}(-z_{i},\hat{\theta}_{t}),\\ \mathcal{I}_{\textrm{IP Ensemble}}^{\textrm{util}}(-z_{i},\Theta_{\textrm{IP Ensemble}})&=\frac{1}{T}\sum\nolimits_{\hat{\theta}_{t}\in\Theta_{\textrm{IP Ensemble}}}\mathcal{I}^{\textrm{util}}_{\textup{IP}}(-z_{i},\hat{\theta}_{t}),\\ \end{split}$ (5)
where $\hat{\theta}$ in IP is the converged model parameters; $\Theta_{\textrm{TracIn}}$ denotes the set of model parameters from the saved checkpoints during model training; $\Theta_{\textrm{GEX}}$ is obtained by training the converged model for several extra iterations; IP Ensemble gets $\Theta_{\textrm{IP Ensemble}}$ by applying dropout mechanisms on the model; $T$ is the ensemble size. The advantages of IP Ensemble are that it does not require recording checkpoints compared to TracIn, and it does not need extra model training processes compared to GEX. We argue that the early checkpoints might not be helpful to analyze the sample influence and the extra iterations on the converged model make little effects.
Figure 8: Illustrating our inner product approach on measuring fairness and robustness. A and D illustrate the same 2D linearly separable synthetic dataset in Figure 2 trained using a Logistic Regression model for binary classification, where the solid and dashed point boundaries denote the majority and minority subgroups. B and E represent the validation set. C and F show the estimated influence on fairness and robustness by the traditional influence function and our IP.
@@ C.1 IP for measuring fairness and robustness
Similar to Figure 2, we conduct the experiments based on Eqs. (3) and (4) to demonstrate the effectiveness of our IP in measuring the sample influence on fairness and robustness. Figure 8 shows the relationship between IP and the traditional influence function, indicate that IP is a good surrogate of the traditional influence function in measuring model’s fairness and robustness as well.
@@ C.2 Performance of IP Ensemble with different rates of removed samples
We also explore different rates of removed samples on CIFAR-10N-a. Table 4 shows the performance of IP Ensemble with different rates of removed samples. Except for not removing, there is no significant difference in remaining results.
@@ C.3 Parameter Analysis on Dropout Rate and Ensemble Size
C.4 Performance on ViT and MLP-Mixer
To verify the effectiveness of our IP ensemble across different network architectures, we conducte experiments using ViT [15] and MLP-Mixer [54] on various vision datasets in Table 6. We use a batch size of 512 for all experiments in this section. The models are trained on CIFAR-10N and CIFAR-100N datasets for 100 epochs, and on Animal-10N for 400 epochs. The learning rate was set to $1\times 10^{-4}$ for ViT and $1\times 10^{-3}$ for MLP-Mixer.
Our IP Ensemble consistently demonstrates superior performance compared to the vanilla models and other baseline methods on different base models. Specifically, when trained on ViT, IP Ensemble achieves an average accuracy of 70.06%, outperforming the vanilla ViT’s average accuracy of 68.87%. This improvement is observed across all datasets, with notable gains on CIFAR-100N (51.01% versus 48.53%) and Animal-10N (77.20% versus 76.20%). Additionally, compared to baseline methods, our IP Ensemble achieves the best performance with ViT or MLP-Mixer based models, indicating the effectiveness of our methods on different base models.
Table 6: Accuracy results of influence function-based methods with ViT and MLP-mixer as based models on CIFAR10N-a, CIFAR-100N, and Animal-10N with 5% detrimental samples removed.
Methods / Datasets CIFAR-10N-a CIFAR-100N Animal-10N Avg.
ViT 81.87 48.53 76.20 68.87
LiSSA [26] 81.83 $\pm$ 0.19 48.47 $\pm$ 0.24 76.90 $\pm$ 0.35 69.07
TracIn [46] 81.59 $\pm$ 0.21 48.18 $\pm$ 0.28 76.21 $\pm$ 0.58 68.66
EKFAC [19] 80.79 $\pm$ 0.56 48.90 $\pm$ 0.30 76.85 $\pm$ 0.30 68.85
DataInf [31] 81.68 $\pm$ 0.21 49.00 $\pm$ 0.22 76.70 $\pm$ 0.27 69.13
Self-TracIn [52] 81.88 $\pm$ 0.12 49.25 $\pm$ 0.21 76.95 $\pm$ 0.25 69.36
Self-LiSSA [5] 82.89 $\pm$ 0.12 49.30 $\pm$ 0.18 76.80 $\pm$ 0.28 69.66
TDA [3] 81.89 $\pm$ 0.28 49.05 $\pm$ 0.54 76.44 $\pm$ 0.32 69.12
GEX [25] 82.99 $\pm$ 0.24 49.78 $\pm$ 0.20 76.88 $\pm$ 0.29 69.88
IP (Ours) 81.93 $\pm$ 0.22 49.12 $\pm$ 0.20 76.80 $\pm$ 0.22 69.28
IP Ensemble (Ours) 81.96 $\pm$ 0.13 51.01 $\pm$ 0.17 77.20 $\pm$ 0.24 70.06
MLP-Mixer 74.42 35.63 72.69 60.91
LiSSA [26] 75.35 $\pm$ 0.47 36.50 $\pm$ 0.21 72.89 $\pm$ 0.28 61.58
TracIn [46] 74.43 $\pm$ 0.28 35.21 $\pm$ 0.35 72.71 $\pm$ 0.43 60.78
EKFAC [19] 75.49 $\pm$ 0.28 36.34 $\pm$ 0.24 73.12 $\pm$ 0.19 61.65
DataInf [31] 75.10 $\pm$ 0.45 35.78 $\pm$ 0.29 73.17 $\pm$ 0.21 61.35
Self-TracIn [52] 75.88 $\pm$ 0.21 36.72 $\pm$ 0.19 73.81 $\pm$ 0.54 62.14
Self-LiSSA [5] 75.41 $\pm$ 0.38 36.05 $\pm$ 0.27 72.56 $\pm$ 0.34 61.34
TDA [3] 74.89 $\pm$ 0.31 36.84 $\pm$ 0.38 71.88 $\pm$ 0.43 61.20
GEX [25] 75.57 $\pm$ 0.27 36.85 $\pm$ 0.25 73.89 $\pm$ 0.43 62.10
IP (Ours) 75.04 $\pm$ 0.23 36.15 $\pm$ 0.18 73.47 $\pm$ 0.20 61.55
IP Ensemble (Ours) 75.77 $\pm$ 0.19 37.12 $\pm$ 0.18 73.80 $\pm$ 0.11 62.23
C.5 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Table 7 shows the time complexity of various influence-based methods. Except for the vanilla one, all other methods have linear time complexity in terms of the sample size. However, they have large divergence in real execution time.
Table 7: Computational complexity of influence-function-based methods ($n$ is #training samples and $p$ is #model parameters, $k$ is #checkpoints or #ensemble size with $k$$=$$1$ here. "-" denotes no runs.)
@@ C.5 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Self-TracIn [52] Self-influence $\mathcal{O}(npk)$ 0.15 1.41 0.15
@@
GEX [25] Hessian-free $\mathcal{O}(npk)$ 0.03 0.28 0.03
TDA [3] Hessian-based $\mathcal{O}(np^{2}k)$ 23.98 193.79 24.77
@@ Appendix D Broader Impact and Limitations
Our work aims to address issues that currently hinder the applicability of influence estimation in deep learning models. By enabling influence estimation for deep models, practitioners can assess whether training samples are beneficial or detrimental to performance. As we demonstrate through extensive experiments across multiple problem settings, our proposed outlier gradient analysis approach outperforms existing baselines and can augment model performance by trimming detrimental samples. As a result, our work paves the way for significant positive societal impact, especially with the increased adoption of larger and deeper neural networks such as LLMs.
However, despite our research demonstrating superior performance and potential benefits, there are several limitations to consider. First, our method assumes a level of model and data homogeneity that might not hold in highly heterogeneous datasets or more complex model architectures. Furthermore, there is the consideration of robustness and fairness across different domains and types of data; our current evaluations, while extensive, may not cover all possible scenarios or edge cases, potentially limiting the generalizability of our findings. Finally, ethical considerations around data trimming and sample selection must be carefully managed to avoid unintended biases or negative impacts on model performance for underrepresented groups. As influence estimation becomes more integrated into model training and evaluation pipelines, ongoing research will be essential to ensure that these techniques are applied responsibly and equitably.
Due to the extensive computational demands of our experiments (see details in Appendix E), which require multiple retraining cycles to evaluate model stability and robustness, it becomes impractical for us to conduct comparative experiments on very large datasets, e.g., ImageNet. The limited computational resources at our disposal make it challenging to perform such large-scale retraining efforts, especially given the significant training time and GPU requirements. As a result, we focus our evaluation on middle-size yet representative benchmarks that allow for more feasible experimentation while maintaining the rigor of our comparative analyses.